Units produced = 120,000; Standard Direct labor hours per Unit = 0.35; Standard variable overhead rate per direct labor hour = $4; Actual variable overhead costs = $105,000; Actual hours worked = 30,000.

Calculate variable overhead spending variance and variable overhead efficiency variance
Answer #1

Variable overhead spending variance

= Standard voh - Actual voh

= (120,000*0.35*4) - 105,000

= 63,000 Favourable

.

Voh efficiency variance = (SLH - ALH) * SVOHR

= (120,000*0.35 - 30,000) * 4

= 48,000 Favourable